# biblio.bib

### v2

@@ Abstract
We consider the logarithmic Fekete problem, which consists of placing a fixed number of points on the unit sphere in $\mathbb{R}^{d}$, in such a way that the product of all pairs of mutual Euclidean distances is maximized or, equivalently, so that their logarithmic energy is minimized. Using tools from Computational Algebraic Geometry, we find and classify all critical configurations for this problem when considering at most six points in every dimension $d$. In particular, our approach gives new proofs of several key results appearing in the literature, with the benefit of using a unified approach.
@@ 1 Introduction
This is considered a highly non-trival optimization problem, with exact solutions known only for a few values of $n$. Indeed, Smale’s 7th problem, listed among the key open problems for the 21st century by Steve Smale, asks whether it is possible to find $n$ points on the sphere $S^{2}$ in polynomial time in $n$, so that its logarithmic energy differs from the optimal value by at most $c\ln n$, for a universal constant $c$ [smale_problems]. A key difficulty of this problem is that the optimal value remains unknown to logarithmic precision [betermin2018renormalized, carlos_fatima].
There are several lines of research related to this problem. Among them, we highlight two. One focuses on constructing random or deterministic configurations with “good” energy values for an arbitrary number of points, usually large or asymptotic [ABS, alishahi2015spherical, beltran2018diamond]. The other aims to characterize optimal configurations for specific values of $n$, typically small ones. Our work falls within this latter setting.
@@ 2 Related work
For $S^{1}$, the solution of Problem (2) is known for any number of points, and is given by $n$ equidistributed points [borodachov2019, Theorem 2.3.3]. We denote this configuration as $n$-gon or “Equator”. For $S^{2}$, the solutions are known only for up to six points [andreev1996, kolushov1997, dragnev2002], and for twelve points [andreev1996]. For $S^{3}\subset\mathbb{R}^{4}$ and six points, the solution is given by two equilateral triangles, inscribed in two great circles $S^{1}$, orthogonal to each other [Dragnev2016, Theorem 1.9]. We denote this configuration as $3_{S^{1}}\times 3_{S^{1}}$. Table 1 lists the known solutions for up to six points, for all possible dimensions of the sphere. We use the term $(n-1)$-simplex for the configuration that has $n$ points on the sphere $S^{d-1}$, all at the same distance from each other. The $(n-1)$-simplex is the optimal configuration for $n$ points on $S^{d-1}$, whenever $d+1\geq n\geq 3$ [kolushov1997]. Note that the 3-simplex is the regular Tetrahedron.
@@
4 4-gon Tetrahedron [kolushov1997] N/A N/A
5 5-gon Bipyramid [dragnev2002] 4-simplex [kolushov1997] N/A
6 6-gon Octahedron [kolushov1997] $3_{S^{1}}\times 3_{S^{1}}$ [Dragnev2016] 5-simplex [kolushov1997]
Regarding critical configurations of Problem (2), little is known about them, even for the cases where the solutions are known. One thing that is known is that the problem has no local maxima [beltran2013, Corollary 1.3], and that critical configurations always have center of mass zero [dragnev2002, Proposition 2]. It is also known that the global minima is not always unique. In particular, if $q=p^{l}$, with $p>2$ prime and $l\geq 1$, taking $n=(q+1)(q^{3}+1)$ points on $S^{d-1}$, $d=q\frac{q^{3}+1}{q+1}$, there are $\lfloor(l-1)/2\rfloor$ essentially different global minima [ballinger2009, Section 1]. On the other hand, in [Dragnev2016, Dragnev2023], the authors study the critical configurations for $d+2$ points on $S^{d-1}$, and give a method to classify all critical configurations, and to characterize the non-degenerate ones. Moreover, for six points in $S^{3}$, they find the global minimum and give the first known case of a spurious local minima. Also, the only saddle points that are explicitly mentioned in the literature are the Equator [shub1993, Section 2], and two configurations for six points in $S^{3}$ [Dragnev2016, Dragnev2023]. Another natural question is whether every saddle may be classified using the Hessian. This is particularly important for numerical optimization [ge2015, jin2017]. For 7 points on $S^{2}$, there is a critical configuration that the Hessian does not classify [constantineau2023, Remark 6.3], although it is believed that it is a global minimum. Finally, based on numerical experiments, it is conjectured that the number of spurious local minima in $S^{2}$ increases “dramatically” with the number of points [rakhmanov1995, Section 4], although no spurious local minima are known in this dimension.
@@
Although the optimal configuration for six points in $S^{2}$ was already reported, our method obtains the first exhaustive list of all critical configurations and their classification.
For six points in $S^{3}$, since our method obtains all critical configurations, including the degenerate ones, this complements the results of [Dragnev2016, Dragnev2023], giving an exhaustive list of all critical configurations in this case.
3 System of critical configurations
@@
$\frac{\partial L(w,\lambda)}{\partial w_{k}}=-\sum_{j=1,j\neq k}^{n}\frac{2\left(w_{k}-w_{j}\right)}{\|w_{k}-w_{j}\|^{2}}+2\lambda_{k}w_{k}=\vec{0},\quad\forall\ k=1,\ldots,n.$
@@
$\lambda_{k}=\frac{n-1}{2},\quad\forall\ k=1,\ldots,n.$
@@
$\|w_{k}\|^{2}=1,\quad\forall\ k=1,\ldots,n,$ (3)
@@
$\frac{n-1}{2}w_{k}-\sum_{j=1,j\neq k}^{n}\frac{w_{k}-w_{j}}{\|w_{k}-w_{j}\|^{2}}=\vec{0},\quad\forall\ k=1,\ldots,n.\$ (4)
@@
4 Removing orthogonal symmetry
@@
$\sum_{j=1,j\neq k}^{n}\frac{w_{i}^{T}w_{k}-w_{i}^{T}w_{j}}{\|w_{k}-w_{j}\|^{2}}=\frac{n-1}{2}w_{i}^{T}w_{k},\quad\forall\ k,i=1,\ldots,n.$
As $w_{l}^{T}w_{l}=1,\forall\ l$, we have: $\|w_{k}-w_{j}\|^{2}=2(1-w_{k}^{T}w_{j})$. Using this, the equations may be written in the new variables as:
$\sum_{j=1,j\neq k}^{n}\frac{x_{ik}-x_{ij}}{1-x_{kj}}=(n-1)x_{ik},\quad\forall\ k\neq i.$ (6)
@@
$\sum_{k=1}^{n}w_{j}^{T}w_{k}=0,\quad\forall\ j=1,\ldots,n.$
@@
$1+\sum_{k=1,k\neq j}^{n}x_{jk}=0,\quad\forall\ j=1,\ldots,n.$ (7)
@@
5 Relation between solutions of both systems
@@
[…]
@@
Proposition 1 (Autonne-Takagi factorization [horn2013, Corollary 2.6.6]).
@@
We call $X$ the dot product matrix, as we may write $X=W^{T}W$. Thus, given a solution $X\in\mathbb{C}^{n\times n}$ of rank $d$, we obtain a configuration $(w_{1},...,w_{n})$ given by the columns of $W\in\mathbb{C}^{d\times n}$. As the relation between both equation systems is obtained by taking dot product with $w_{i}$, $i=1,...,n$, and this set is a generator of $\mathbb{C}^{d}$, we get that both systems are equivalent. Thus, we do not lose or introduce solutions in $\mathbb{C}$. The next result characterizes the correspondence between real solutions of both systems, which is the ones we are really interested in.
Proposition 2 ([horn2013, Corollary 2.5.11]).
A matrix $X\in\mathbb{R}^{n\times n}$ is symmetric and positive semi-definite if and only if there exists $W\in\mathbb{R}^{d\times n}$, such that $X=W^{T}W$. In this case, if $X=Q^{T}DQ$ is an orthogonal eigenvalue factorization of $X$, we can take $W:=\sqrt{D}Q$.
@@
6 Polynomial system of equations
@@
$z_{kj}\left(1-x_{kj}\right)=1,\quad\forall\ k\neq j.$ (8)
@@
$\sum_{j=1,j\neq k}^{n}\left(x_{ik}-x_{ij}\right)z_{kj}=(n-1)x_{ik},\quad\forall\ k\neq i.$ (9)
@@
equations $n(3n-1)/2$ 12 22 35 51 70 92
7 Counting complex solutions
To apply our method, we need to verify that the system of polynomial equations (7), (8) and (9) has a finite number of solutions, and to count this number. We first define the polynomial ideal $I\subset\mathbb{Q}[z_{ij},x_{ij}]$, generated by the polynomials of the equations. The set of complex solutions of these equations is denoted by $V(I)\subset\mathbb{C}^{m}$, where $m$ is the number of variables. We then use the following result, which allows us to compute an upper bound on the true number of critical points via Gröbner bases.
Theorem 1 ([cox2005, Finiteness Th. p. 39, Corollary 2.5 Ch. 4]).
Let $I$ be an ideal of the ring $R=\mathbb{C}[x_{1},\ldots,x_{m}]$, and $G$ a Gröbner basis of $I$, w.r.t. a graded monomial order. The set of solutions $V(I)\subset\mathbb{C}^{m}$ is finite, if and only if, for every variable $x_{i}$, there exists $\alpha_{i}\in\mathbb{Z}^{+}$, such that $x_{i}^{\alpha_{i}}\in LT(G)$; where $LT(G)$ denotes the set of leading terms of the elements of $G$. In this case, the number of complex solutions of $V(I)$, counted with multiplicity, is the number of monomials of $R$ that are not in the ideal $\left\langle LT(G)\right\rangle$ (not divisible by any element of $LT(G)$).
To calculate a Gröbner basis we use msolve [berthomieu2021], with graded reverse lexicographic monomial order (grevlex). We then pass this basis to Macaulay2 (M2) [M2] and use it to verify that the number of solutions is finite and to obtain the number of solutions $(\deg I)$. Both msolve and M2 are open source software. Msolve implements the F4 algorithm [faugere1999], with parallel algorithms for solving the linear systems generated by F4. Table 3 shows the “expected number of solutions” of the polynomial system for different number of points. This is equal to the exact number of (possibly complex) solutions, counted with multiplicity, as defined in Section 1.
Table 3: Expected solutions of polynomial system for $n\leq 6$. Time and RAM of msolve to calculate Gröbner basis (grevlex).
@@ 7 Counting complex solutions
8 Counting permutations
@@
9 Energies and optimal configurations
@@
10 Four and five points
@@
$X_{\text{Tet.}}=\begin{pmatrix}1&\theta&\theta&\theta\\ \theta&1&\theta&\theta\\ \theta&\theta&1&\theta\\ \theta&\theta&\theta&1\\ \end{pmatrix},\quad\theta:=-\frac{1}{3}.$
@@
Conf. $|Orb(X)|$ Energy $E/2^{15}$ $rank(X)$
@@
Five points. We have at least three candidates for critical configurations in $S^{2}$: the Equator and two other denoted 1:3:1 and 1:4. (Föppl notation [foppl1912]). Configuration 1:3:1 has two points forming a dipole, and three other points equidistributed in a plane through the origin, orthogonal to the dipole. The dot product matrix is:
$X_{1:3:1}=\begin{pmatrix}1&A&0&0&0\\ A&1&0&0&0\\ 0&0&1&B&B\\ 0&0&B&1&B\\ 0&0&B&B&1\end{pmatrix},\quad\begin{array}[]{l}A:=-1,\\ B:=-\frac{1}{2}\end{array}.$
@@
$X_{1:4}=\begin{pmatrix}1&A&A&A&A\\ A&1&B&C&B\\ A&B&1&B&C\\ A&C&B&1&B\\ A&B&C&B&1\\ \end{pmatrix},\quad\begin{array}[]{l}A:=-\frac{1}{4},\\ B:=\frac{1}{16},\\ C:=-\frac{7}{8}\end{array}.$
@@
$X_{\text{Eq.}}=\begin{pmatrix}1&A&B&B&A\\ A&1&A&B&B\\ B&A&1&A&B\\ B&B&A&1&A\\ A&B&B&A&1\end{pmatrix},\quad\begin{array}[]{l}A:=\frac{-1+\sqrt{5}}{4},\\ B:=\frac{-1-\sqrt{5}}{4}\end{array}.$
@@
Configuration $|Orb(X)|$ Energy $E/2^{15}$ $rank(X)$

### v3

# Characterization of Logarithmic Fekete Critical Configurations of at Most Six Points in All Dimensions

@@ Abstract
We consider the logarithmic Fekete problem, which consists of placing a fixed number of points on the unit sphere in $\mathbb{R}^{d}$, in such a way that the product of all pairs of mutual Euclidean distances is maximized or, equivalently, so that their logarithmic energy is minimized. Using tools from Computational Algebraic Geometry, we find and classify all critical configurations for this problem when considering at most six points in every dimension $d$. In particular, our approach gives new proofs of several key results appearing in the literature, with the benefit of using a unified approach. Furthermore, for seven points in $S^{2}$, we characterize the global minimizer among critical configurations having at least one pair of antipodal points, and give numerical evidence to support the conjecture that this configuration is also the unrestricted global minimizer.
Keywords: Fekete, logarithmic energy, critical points, Gröbner basis, msolve.
@@ 1 Introduction
This is considered a highly non-trivial optimization problem, with exact solutions known only for a few values of $n$. Indeed, Smale’s 7th problem, listed among the key open problems for the 21st century by Steve Smale, asks whether it is possible to find $n$ points on the sphere $S^{2}$ in polynomial time in $n$, so that its logarithmic energy differs from the optimal value by at most $c\ln n$, for a universal constant $c$ [29]. A key difficulty of this problem is that the optimal value remains unknown to logarithmic precision [10, 6].
There are several lines of research related to this problem. Among them, we highlight two. One focuses on constructing random or deterministic configurations with “good” energy values for an arbitrary number of points, usually large or asymptotic [3, 1, 5]. The other aims to characterize optimal configurations for specific values of $n$, typically small ones. Our work falls within this latter setting.
@@ 2 Related work
For $S^{1}$, the solution of Problem (2) is known for any number of points, and is given by $n$ equidistributed points [11, Theorem 2.3.3]. We denote this configuration as $n$-gon or “Equator”. For $S^{2}$, the solutions are known only for up to six points [2, 26, 15], and for twelve points [2]. For $S^{3}\subset\mathbb{R}^{4}$ and six points, the solution is given by two equilateral triangles, inscribed in two great circles $S^{1}$, orthogonal to each other [17, Theorem 1.9]. We denote this configuration as $3_{S^{1}}\times 3_{S^{1}}$. Table 1 lists the known solutions for up to six points, for all possible dimensions of the sphere. We use the term (n1)(n-1)-simplex for the configuration that has $n$ points on the sphere $S^{d-1}$, all at the same distance from each other. The (n1)(n-1)-simplex is the optimal configuration for $n$ points on $S^{d-1}$, whenever $d+1\geq n\geq 3$ [26]. Note that the 3-simplex is the regular Tetrahedron.
@@
4 4-gon Tetrahedron [26] N/A N/A
5 5-gon Bipyramid [15] 4-simplex [26] N/A
6 6-gon Octahedron [26] $3_{S^{1}}\times 3_{S^{1}}$ [17] 5-simplex [26]
Regarding critical configurations of Problem (2), little is known about them, even for the cases where the solutions are known. One thing that is known is that the problem has no local maxima [7, Corollary 1.3], and that critical configurations always have center of mass zero [15, Proposition 2]. It is also known that the global minima is not always unique. In particular, if $q=p^{l}$, with $p>2$ prime and $l\geq 1$, taking $n=(q+1)(q^{3}+1)$ points on $S^{d-1}$, $d=q\frac{q^{3}+1}{q+1}$, there are $\lfloor(l-1)/2\rfloor$ essentially different global minima [4, Section 1]. On the other hand, in [17, 16], the authors study the critical configurations for $d+2$ points on $S^{d-1}$, and give a method to classify all critical configurations, and to characterize the non-degenerate ones. Moreover, for six points in $S^{3}$, they find the global minimum and give the first known case of a spurious local minima. Also, the only saddle points that are explicitly mentioned in the literature are the Equator [28, Section 2], and two configurations for six points in $S^{3}$ [17, 16]. Another natural question is whether every saddle may be classified using the Hessian. This is particularly important for numerical optimization [21, 25]. For 7 points on $S^{2}$, there is a critical configuration that the Hessian does not classify [13, Remark 6.3], although it is believed that it is a global minimum [8][Conjecture 11.1]. Finally, based on numerical experiments, it is conjectured that the number of spurious local minima in $S^{2}$ increases “dramatically” with the number of points [27, Section 4], although no spurious local minima are known in this dimension.
@@
Although the optimal configuration for six points in $S^{2}$ was already reported, our method obtains the first exhaustive list of all critical configurations and their classification. For six points in $S^{3}$, since our method obtains all critical configurations, including the degenerate ones, this complements the results of [17, 16], giving an exhaustive list of all critical configurations in this case.
Finally, for seven points in $S^{2}$, we prove that if a solution to the Fekete problem has a dipole, then it is given by configuration 1:5:1, which is the solution conjectured via numerical experiments. If 1:5:1 is indeed the global minimum, our theorem reduces the problem of proving that it is optimal to verifying that some minimizer of the Fekete problem must have a dipole. We also provide numerical evidence in favor of the optimality of 1:5:1.
3 Formulation and computational algebraic tools
3.1 System of critical configurations
@@
$\frac{\partial L(w,\lambda)}{\partial w_{k}}=-\sum_{j=1,j\neq k}^{n}\frac{2\left(w_{k}-w_{j}\right)}{\|w_{k}-w_{j}\|^{2}}+2\lambda_{k}w_{k}=\vec{0},\qquad\forall\ k=1,\ldots,n.$
@@
$\lambda_{k}=\frac{n-1}{2},\qquad\forall\ k=1,\ldots,n.$
@@
$\|w_{k}\|^{2}=1,\qquad\forall\ k=1,\ldots,n,$ (3)
@@
$\frac{n-1}{2}w_{k}-\sum_{j=1,j\neq k}^{n}\frac{w_{k}-w_{j}}{\|w_{k}-w_{j}\|^{2}}=\vec{0},\qquad\forall\ k=1,\ldots,n.\\$ (4)
@@
3.2 Removing orthogonal symmetry
@@
$\sum_{j=1,j\neq k}^{n}\frac{w_{i}^{T}w_{k}-w_{i}^{T}w_{j}}{\|w_{k}-w_{j}\|^{2}}=\frac{n-1}{2}w_{i}^{T}w_{k},\qquad\forall\ k,i=1,\ldots,n.$
As $w_{l}^{T}w_{l}=1,\forall\ l=1,\ldots,n$, we have: $\|w_{k}-w_{j}\|^{2}=2(1-w_{k}^{T}w_{j})$. Using this, the equations may be written in the new variables as:
$\sum_{j=1,j\neq k}^{n}\frac{x_{ik}-x_{ij}}{1-x_{kj}}=(n-1)x_{ik},\qquad\forall\ k\neq i.$ (6)
@@
$\sum_{k=1}^{n}w_{j}^{T}w_{k}=0,\qquad\forall\ j=1,\ldots,n.$
@@
$1+\sum_{k=1,k\neq j}^{n}x_{jk}=0,\qquad\forall\ j=1,\ldots,n.$ (7)
@@
3.3 Relation between solutions of both systems
@@
Let $x_{ij}\in\mathbb{C},\ 1\leq i<j\leq n$. Define $X\in\mathbb{C}^{n\times n}$, symmetric, with $X_{ij}=x_{ij},\ \forall\ i<j$, $X_{ii}=1,\ \forall\ i$.
@@
Proposition 1 (Autonne-Takagi factorization [24, Corollary 2.6.6]).
@@
We call $X$ the dot product matrix, as we may write $X=W^{T}W$. Thus, given a solution $X\in\mathbb{C}^{n\times n}$ of rank $d$, we obtain a configuration $(w_{1},...,w_{n})$ given by the columns of $W\in\mathbb{C}^{d\times n}$. As the relation between both equation systems is obtained by taking dot product with $w_{i}$, i=1,,ni=1,...,n, and this set is a generator of $\mathbb{C}^{d}$, we get that both systems are equivalent. Thus, we do not lose or introduce solutions in $\mathbb{C}$. The next result characterizes the correspondence between real solutions of both systems, which is the ones we are really interested in.
Proposition 2 ([24, Corollary 2.5.11]).
A matrix $X\in\mathbb{R}^{n\times n}$ is symmetric and positive semi-definite if and only if there exists $W\in\mathbb{R}^{d\times n}$, such that $X=W^{T}W$, for some $d\in\mathbb{N}$. In this case, if $X=Q^{T}DQ$ is an orthogonal eigenvalue factorization of $X$, we can take $W:=\sqrt{D}Q$.
@@
3.4 System of polynomial equations
@@
$z_{kj}\left(1-x_{kj}\right)=1,\qquad\forall\ k\neq j.$ (8)
@@
$\sum_{j=1,j\neq k}^{n}\left(x_{ik}-x_{ij}\right)z_{kj}=(n-1)x_{ik},\qquad\forall\ k\neq i.$ (9)
@@
equations n(3n1)/2n(3n-1)/2 12 22 35 51 70 92
Note that this formulation gives the critical configurations for all sphere dimensions, and not just for $S^{2}$. However, it is possible to add polynomial constraints to select critical configurations only up to a given sphere dimension $S^{k-1}\subset\mathbb{R}^{k}$. For example, by adding all $(k+1)\times(k+1)$ minors of the $n\times n$ dot product matrix $X$, which implies $\text{rank}(X)\leq k$. This adds ${n\choose k+1}\times{n\choose k+1}$ equations to the polynomial system, each of degree $k+1$. Table 3 shows the number of equations that should be added for different values of $n$ and $k\leq n-1$. We will not include these constraints, since we are interested in characterizing the critical configurations in all dimensions.
Table 3: Number of additional constraints to impose $\text{rank}(X)\leq k$ using all $(k+1)\times(k+1)$ minors of $X$.
[…]
$n$ 2 3 4 5
5 25 1
6 225 36 1
7 1225 441 49 1
3.5 Counting complex solutions
To apply our method, we need to verify that the system of polynomial equations (7), (8) and (9) has a finite number of solutions, and to count this number. We first define the ideal $I\subset\mathbb{Q}[z_{ij},x_{ij}]$, generated by the equations. The set of complex solutions of these equations is denoted by $V(I)\subset\mathbb{C}^{m}$, where $m$ is the number of variables. We then use the following result, which allows us to compute an upper bound on the true number of critical points via Gröbner bases.
Theorem 1 ([14, Finiteness Th. p. 39, Corollary 2.5 Ch. 4]).
Let $I\subset\mathbb{Q}[x_{1},\ldots,x_{m}]$ be an ideal, and $G$ a Gröbner basis of $I$. The set of complex solutions $V(I)\subset\mathbb{C}^{m}$ is finite, if and only if, for every variable $x_{i}$, there exists an integer $\alpha_{i}\geq 0$, such that $x_{i}^{\alpha_{i}}\in LM(G)$; where $LM(G)$ denotes the set of leading monomials of the elements of $G$. In this case, the number of complex solutions, counted with multiplicity, is the number of monomials of the ring that are not in the ideal $\left\langle LM(G)\right\rangle$ (not divisible by any element of $LM(G)$).
To calculate a Gröbner basis we use msolve [9], with graded reverse lexicographic monomial order (grevlex). We then pass this basis to Macaulay2 (M2) [22] and use it to verify that the number of solutions is finite and to obtain the number of solutions $(\deg I)$. Both msolve and M2 are open source software. Msolve implements the F4 algorithm [19], with parallel algorithms for solving the linear systems generated by F4. Table 4 shows the “expected number of solutions” of the polynomial system for different number of points. This is equal to the exact number of (possibly complex) solutions, counted with multiplicity, as defined in Section 1.
Table 4: Expected solutions of polynomial system for $n\leq 6$. Time and RAM of msolve v.0.73 to calculate a Gröbner basis (grevlex) using 20 threads.
@@ 3.5 Counting complex solutions
Note that, for counting the number of solutions, we only need the leading monomials of a Gröbner basis. We opt to calculate the whole basis, and in particular to lift all the basis coefficients to the rationals, to inform the resources needed, as this basis could be used for other purposes in the stage of finding solutions.
3.6 Counting permutations
@@
3.7 Energies and optimal configurations
@@
4 Results for $n\leq 7$ points
In what follows, we apply our method to the different formulations presented above and for different numbers of points. This allows us to recover some known results and to prove new theorems in previously unexplored cases.
4.1 Four and five points
@@
$X_{\text{Tet.}}=\begin{pmatrix}1&\theta&\theta&\theta\\ \theta&1&\theta&\theta\\ \theta&\theta&1&\theta\\ \theta&\theta&\theta&1\\ \end{pmatrix},\qquad\theta:=-\frac{1}{3}.$
@@
Conf. $|Orb(X)|$ Energy $E/2^{n\choose 2}$ $rank(X)$
@@
Five points. We have at least three candidates for critical configurations in $S^{2}$: the Equator and two other denoted 1:3:1 and 1:4. (Föppl notation [20]). Configuration 1:3:1 has two points forming a dipole, and three other points equidistributed in a plane through the origin, orthogonal to the dipole. The dot product matrix is:
$X_{1:3:1}=\begin{pmatrix}1&A&0&0&0\\ A&1&0&0&0\\ 0&0&1&B&B\\ 0&0&B&1&B\\ 0&0&B&B&1\end{pmatrix},\qquad\begin{array}[]{l}A:=-1,\\ B:=-\frac{1}{2}\end{array}.$
@@
$X_{1:4}=\begin{pmatrix}1&A&A&A&A\\ A&1&B&C&B\\ A&B&1&B&C\\ A&C&B&1&B\\ A&B&C&B&1\\ \end{pmatrix},\qquad\begin{array}[]{l}A:=-\frac{1}{4},\\ B:=\frac{1}{16},\\ C:=-\frac{7}{8}\end{array}.$
@@
$X_{\text{Eq.}}=\begin{pmatrix}1&A&B&B&A\\ A&1&A&B&B\\ B&A&1&A&B\\ B&B&A&1&A\\ A&B&B&A&1\end{pmatrix},\qquad\begin{array}[]{l}A:=\frac{-1+\sqrt{5}}{4},\\ B:=\frac{-1-\sqrt{5}}{4}\end{array}.$
@@
Configuration $|Orb(X)|$ Energy $E/2^{n\choose 2}$ $rank(X)$